OpenStudy (anonymous):

.a cylinder volume is 96 cubic feet. what is the volume of the cone

OpenStudy (triciaal):

WELCOME

OpenStudy (triciaal):

volume of a cylinder = pi*r^2* h r is the radius and h is the height volume of a cone = 1/3*pi*r^2*h

OpenStudy (triciaal):

if the cone fits in the cylinder then the radius and height will be the same the the volume of the cone will be 1/3 the volume of the cylinder if this is true then volume of the cone = 1/3*96 = 32

OpenStudy (triciaal):

unit of measure for volume here = cubic feet